The radius of the outer spear of a spherical capacitor is five times the radius of its inner shell . What are the dimensions of these capacitor if its capacitance is 5.00 picofarad?


1
Expert's answer
2021-05-14T09:42:31-0400

Radius of the outer shell: R=5r

Radius of the inner shell: r


Solution

The capacitance is


"C=\frac{4\pi\epsilon_0}{1/r-1/(5r)}=5\pi\epsilon_0r,\\\space\\
r=\frac{C}{5\pi\epsilon_0}=0.036\text{ m},\\\space\\
R=0.18\text{ m}."